Find an equation in the form y=ax^2+bx+c for the parabola passing through the points. (3,-60), (-1,0), (5,-138)

Answer by TimothyLamb(4379)

y = axx + bx + c
---
(3,-60):
9a + 3b + c = -60
---
(-1,0):
a - b + c = 0
---
(5,-138):
25a + 5b + c = -138
---
9a + 3b + c = -60
a - b + c = 0
25a + 5b + c = -138
---
put the system of linear equations into standard form
---
substitute:
x for a
y for b
z for c
---
9x + 3y + z = -60
x - y + z = 0
25x + 5y + z = -138
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -4
y = -7
z = -3
---
substitute:
a for x
b for y
c for z
---
solution:
a = -4
b = -7
c = -3
---
answer:
y = axx + bx + c
y = -4xx - 7x - 3
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---